A model of a tree is made using a scale of 1 inch : 25 feet. What is the height of an actual tree if the height of the model is 4 3/8 inches tall?
Answer by checkley77(9412)

1/25=(35/8)/x
x=25*35/8
x=875/8
x=109.375 ft. tall.